QUASI-JAHN-TELLER EFFECT IN CHEMICAL REACTIONS

LICHANG WANG , Department of Chemistry, The Ohio State University, Columbus, OH 43210.

It will be demonstrated that the quasi-Jahn-Teller effect in chemical reactions is a nuclear quantum effect, T^{q}, induced by electronic nonadiabaticity^{a}. Further, T^{q} is equivalent to the Diagonal Born-Oppenheimer Correction. Results including and without T^{q} for a model system will be presented^{b}. The research shows clearly that one has to consider this effect in order to obtain various quantities within a spectroscopic accuracy for systems where this effect is no longer negligible.

aL. Wang, Chem. Phys. Lett. (a) 285 (1998) 359; (b) submitted.

bL. Wang, Chem. Phys. 237 (1998) 305.